# Runge–Kutta Physics Informed Neural Networks: Formulation and Analysis

### v1

@@ 1.1 Evolutionary PDEs and Neural Network Discretizations
Section 5 is dedicated to numerical experiments. In addition to parabolic equations, we also consider the wave equation. While the convergence theory for hyperbolic equations remains to be developed, the formulation of the method can be directly applied by expressing the equation as a first-order evolution system in time.
@@ 2.2 A linear evolution PDE
{$&u_{t}+Lu=f\qquad&&\text{in }&&\varOmega_{T},\\ &u=0\qquad&&\text{on }&&\partial\varOmega\times(0,T],\\ &u=u^{0}\qquad&&\textrm{in }&&\varOmega,\end{aligned}\right.$ (1)
where $f\in L^{2}(\varOmega_{T}),\;u^{0}\in H^{1}_{0}(\varOmega)$, and $L$ is a coercive, self-adjoint, second order elliptic operator. The associated energies used will be the $L^{2}$-residuals
$\mathcal{E}(v)=\int_{0}^{T}\int_{\varOmega}|v_{t}+Lv-f|^{2}{\mathrm{d}}x\,{\mathrm{d}}t+\,|v-u^{0}|_{H^{1}}^{2}$ (2)
defined over smooth enough functions and domains $\varOmega.$ The choice of the $H^{1}$ seminorm for the initial condition, $|v-u^{0}|_{H^{1}},$ is done in order to obtain a balanced energy which is particularly convenient in the analysis; obviously other choices are possible.
@@
$V_{\mathcal{N},0}=\{u_{\theta}:\varOmega\to{\mathbb{R}},\ \text{where }u_{\theta}(x,t)=\varPhi(x)\mathcal{C}_{L}(x,t),\ \text{for some }\mathcal{C}_{L}\in\mathcal{N}\,\}.$ (6)
@@
possesses a solution $v^{\star}\in V_{\mathcal{N}}.$ The integrals appearing in the loss functional ${\mathcal{E}}$ require further discretization to result in a computable loss. As the set $V_{\mathcal{N}}$ is nonlinear, the problem needs to be considered as minimization in the parameter space over ${\mathbb{R}}^{\dim{\mathcal{N}}}$
@@ 2.3 Simple time-discrete Training
To fix notation, let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ define a partition of $[0,T]$ and $J_{n}:=(t_{n},t_{n+1}],$ $k_{n}:=t_{n+1}-t_{n}.$ We denote by $v^{m}(\cdot)$ and $f^{m}(\cdot)$ the values $v(\cdot,t_{m})$ and $f(\cdot,t_{m}).$ Then, we consider the discrete in time quadrature
@@
$\mathcal{G}_{k,IE}(v)=\sum_{n=0}^{N-1}\,k_{n}\,\int_{\varOmega}\big{|}\frac{v_{n+1}-v_{n}}{k_{n}}+Lv_{n+1}-f(t_{n+1})\big{|}^{2}\,\,{\mathrm{d}}x+\,|v-u^{0}|_{H^{1}(\varOmega)}^{2}\,.$ (14)
@@
$\mathcal{G}_{k,EE}(v)\sum_{n=0}^{N-1}\,k_{n}\,\int_{\Omega}\big{|}\frac{v_{n+1}-v_{n}}{k_{n}}+Lv_{n}-f(t_{n})\big{|}^{2}\,\,{\mathrm{d}}x+\,|v-u^{0}|_{H^{1}(\varOmega)}^{2}\,.$ (16)
@@
$\begin{split}\mathcal{G}_{IE,k}(v)&=\|\widehat{v}_{t}+L\overline{v}-\overline{f}\|^{2}_{L^{2}((0,T);L^{2}(\varOmega))}+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}\\ &=\int_{0}^{T}\|\widehat{v}_{t}+L\overline{v}-\overline{f}\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}.\end{split}$ (19)
@@ 2.4.2 Runge–Kutta discrete Loss
We shall introduce more notation related to the representation of piecewise polynomial functions: Let $\ell_{n1},\dotsc,\ell_{nq}\in{\mathbb{P}}_{q-1}$ be the Lagrange polynomials $\ell_{1},\dotsc,\ell_{q}\in{\mathbb{P}}_{q-1}$ for the collocation nodes $c_{1},\dotsc,c_{q}$ shifted to the interval $[t_{n},t_{n+1}],\ell_{ni}(t)=\ell_{ni}(t_{n}+k\tau)=\ell_{i}(\tau),i=1,\dotsc,q.$ Obviously,
@@
for the points $\tilde{c}_{0},\dotsc,\tilde{c}_{q}$ shifted to the interval $[t_{n},t_{n+1}],{\tilde{\ell}}_{ni}(t)={\tilde{\ell}}_{ni}(t_{n}+k\tau)={\tilde{\ell}}_{i}(\tau),i=0,\dotsc,q.$ Obviously, ~${\tilde{\ell}}_{ni}(\tilde{t}_{nj})=\delta_{ij},i,j=0,1,\dotsc,q,$ with $\tilde{t}_{nj}:=t_{n}+k_{n}\tilde{c}_{j},j=0,\dotsc,q.$ With this notation, let the interpolant $\widehat{v}$ of a time dependent function $v$ be the piecewise polynomial function approximating $v$ defined by interpolating the nodal values $v(\tilde{t}_{nj})$ of $v$ as follows
@@
$\begin{split}\mathcal{G}_{RK}(v)&=\|\widehat{v}_{t}(t)+I_{q-1}L\widehat{v}(t)-I_{q-1}f(t)\|^{2}_{L^{2}((0,T);L^{2}(\varOmega))}+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}\\ &=\int_{0}^{T}\|\widehat{v}_{t}(t)+I_{q-1}L\widehat{v}(t)-I_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}.\end{split}$ (26)
@@ 2.4.3 A general discrete Loss
of a time dependent function $v$ to be the piecewise polynomial function of local degree $q$ approximating $v.$ We assume that if $v$ is continuous in time, then $\widehat{v}$ is a piecewise polynomial globally continuous in time function. The standard interpolation operators as defined in Section 2.4.2 are typical cases of these operators. We are ready to define the time-discrete discrete loss by
$\mathcal{G}_{k}(v)=\int_{0}^{T}\|\widehat{v}_{t}(t)+\varPi_{q-1}L\widehat{v}(t)-{\widetilde{\varPi}}_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}.\$ (28)
In Section 4 we show systematically that all Runge-Kutta methods considered and both continuous and discontinuous Galerin time discrete schemes are associated to the general formulation involving the discrete operators $\varPi_{q-1},{\widetilde{\varPi}}_{q-1}$ and ${\widehat{\varPi}}_{q}.$
@@ 3 Stability and Convergence for Parabolic equations
{$&u_{t}+Lu=f\qquad&&\text{in }&&\varOmega_{T},\\ &u=0\qquad&&\text{on }&&\partial\varOmega\times(0,T],\\ &u=u^{0}\qquad&&\textrm{in }&&\varOmega\times\{t=0\}.\end{aligned}\right.$ (30)
@@
$\mathcal{G}(v)=\int_{0}^{T}\|v_{t}(t)+Lv(t)-f(t)\|^{2}_{L^{2}(\varOmega)}{\mathrm{d}}t+|v(0)-u^{0}|^{2}_{H^{1}(\varOmega)}\,.$ (31)
@@ 3.1 Time discrete training through Runge–Kutta and time Galerkin
$\mathcal{G}_{k}(v)=\int_{0}^{T}\|\widehat{v}_{t}(t)+\varPi_{q-1}L\widehat{v}(t)-{\widetilde{\varPi}}_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}^{0}-u^{0}|_{H^{1}(\varOmega)}^{2}.\$
@@
The neural network spaces are selected to meet specific approximability criteria aligned with established results in approximation theory; see, e.g., [Grohs_space_time_approx:2022, Mishra:appr:rough:2022, Xu] and their references. However, subtle challenges arise, particularly with boundary conditions and the fact that existing approximation results do not yet offer concrete guidance on selecting specific architectures. Nevertheless, to investigate the potential convergence of minimizers, we assume that the following approximability requirements are satisfied. The required smoothness of the spaces is guranteed by selecting smooth enough activation functions. The neural network spaces are selected such that for each $\ell\in\mathbb{N}$ we associate a space $V_{\mathcal{N},0},$ which is denoted by $V_{\ell}$ with the approximation property: For each $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega))$ there exists a $w_{\ell}\in V_{\ell}$ such that
@@ 3.1.1 Stability-Maximal Regularity
Following [GGM_pinn_2023], we call our methods stable if the following two key properties, roughly stated as follows, hold
@@
Proof  Since,
$\int_{\varOmega_{T}}(|\widehat{U}_{t}+L\overline{U}|^{2}-2\overline{f}\>(\widehat{U}_{t}+L\overline{U})+|\overline{f}|^{2})\,{\mathrm{d}}x{\mathrm{d}}t\leqslant C,$ (39)
@@ 3.2 Convergence of the minimizers
$\begin{split}\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell},\varphi^{\prime}{\rangle}{\mathrm{d}}t-&\int_{0}^{T}{\langle}\overline{U}_{\ell},\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}\leqslant\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},\varphi^{\prime}-I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}+\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|},\\ &=\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},\varphi^{\prime}-I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}.\end{split}$ (44)
@@
Proof  For $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega)\,)$, we chose a smooth approximation $(w_{\delta})\subset C^{\infty}([0,T];H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega)\,)$ such that
@@
Let $\mathcal{G},\;\mathcal{G}_{\ell}$ be the energy functionals defined in (31) and (14), respectively. Let $u$ be the exact solution of (30) and let $(u_{\ell}),$ $u_{\ell}\in V_{\ell},$ be a sequence of minimizers of $\mathcal{G}_{IE,\ell}$, i.e.,
@@
Applying the Aubin–Lions Lemma, we conclude the existence of $\tilde{u}\in L^{2}((0,T);H^{1}(\varOmega))$ such that $\widehat{u}_{\ell}\rightarrow\tilde{u}$ in $L^{2}((0,T);H^{1}(\Omega))$ up to a subsequence not re-labeled. Furthermore, the arguments in Lemma 3 show that $L\tilde{u}\in L^{2}((0,T);L^{2}(\varOmega))\,.$ Next we show that $\tilde{u}$ is the minimizer of $\mathcal{G},$ and hence $\tilde{u}=u.$ Indeed, let $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega))$, and let $(w_{\ell})$ be such that $w_{\ell}\to w$ and

### v2

@@ 1.1 Evolutionary PDEs and Neural Network Discretizations
Section 5 is dedicated to numerical experiments. In addition to parabolic equations, we also consider the wave equation. While the convergence theory for hyperbolic equations remains to be developed, the formulation of the methods can be directly applied by expressing the equation as a first-order evolution system in time.
@@ 2.2 A linear evolution PDE
{$&u_{t}+Lu=f\qquad&&\text{in }&&\varOmega_{T},\\ &u=0\qquad&&\text{on }&&\partial\varOmega\times(0,T],\\ &u=u_{0}\qquad&&\textrm{in }&&\varOmega,\end{aligned}\right.$ (1)
where $f\in L^{2}(\varOmega_{T}),\;u_{0}\in H^{1}_{0}(\varOmega)$, and $L$ is a coercive, self-adjoint, second order elliptic operator. The associated energies used will be the $L^{2}$-residuals
$\mathcal{E}(v)=\int_{0}^{T}\int_{\varOmega}|v_{t}+Lv-f|^{2}{\mathrm{d}}x\,{\mathrm{d}}t+|v(0)-u_{0}|_{H^{1}}^{2}$ (2)
defined over smooth enough functions and domains $\varOmega.$ The choice of the $H^{1}$ seminorm for the initial condition, $|v(0)-u_{0}|_{H^{1}},$ is done in order to obtain a balanced energy which is particularly convenient in the analysis; obviously, other choices are possible.
@@
$V_{\mathcal{N},0}=\{u_{\theta}:\varOmega_{T}\to{\mathbb{R}},\ \text{where }u_{\theta}(x,t)=\varPhi(x)\mathcal{C}_{L}(x,t),\ \text{for some }\mathcal{C}_{L}\in\mathcal{N}\,\}.$ (6)
@@
possesses a solution $v^{\star}\in V_{\mathcal{N},0}.$ The integrals appearing in the loss functional ${\mathcal{E}}$ require further discretization to result in a computable loss. As the set $V_{\mathcal{N},0}$ is nonlinear, the problem needs to be considered as minimization in the parameter space over ${\mathbb{R}}^{\dim{\mathcal{N}}}$
@@ 2.3 Simple time-discrete Training
To fix notation, let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ define a partition of $[0,T]$ and $J_{n}:=(t_{n},t_{n+1}],$ $k_{n}:=t_{n+1}-t_{n}.$ We denote by $v_{m}(\cdot)$ and $f_{m}(\cdot)$ the values $v(\cdot,t_{m})$ and $f(\cdot,t_{m}).$ Then, we consider the discrete in time quadrature
@@
$\mathcal{G}_{k,IE}(v)=\sum_{n=0}^{N-1}\,k_{n}\,\int_{\varOmega}\big{|}\frac{v_{n+1}-v_{n}}{k_{n}}+Lv_{n+1}-f(t_{n+1})\big{|}^{2}\,\,{\mathrm{d}}x+|v_{0}-u_{0}|_{H^{1}(\varOmega)}^{2}\,.$ (14)
@@
$\mathcal{G}_{k,EE}(v)\sum_{n=0}^{N-1}\,k_{n}\,\int_{\Omega}\big{|}\frac{v_{n+1}-v_{n}}{k_{n}}+Lv_{n}-f(t_{n})\big{|}^{2}\,\,{\mathrm{d}}x+|v_{0}-u_{0}|_{H^{1}(\varOmega)}^{2}\,.$ (16)
@@
$\begin{split}\mathcal{G}_{IE,k}(v)&=\|\widehat{v}_{t}+L\overline{v}-\overline{f}\|^{2}_{L^{2}((0,T);L^{2}(\varOmega))}+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}\\ &=\int_{0}^{T}\|\widehat{v}_{t}+L\overline{v}-\overline{f}\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}.\end{split}$ (19)
@@ 2.4.2 Runge–Kutta discrete Loss
We shall introduce more notation related to the representation of piecewise polynomial functions: Let $\ell_{n1},\dotsc,\ell_{nq}\in{\mathbb{P}}_{q-1}$ be the Lagrange polynomials $\ell_{1},\dotsc,\ell_{q}\in{\mathbb{P}}_{q-1}$ for the collocation nodes $c_{1},\dotsc,c_{q}$ shifted to the interval $[t_{n},t_{n+1}],\ell_{ni}(t)=\ell_{ni}(t_{n}+k_{n}\tau)=\ell_{i}(\tau),i=1,\dotsc,q.$ Obviously,
@@
for the points $\tilde{c}_{0},\dotsc,\tilde{c}_{q}$ shifted to the interval $[t_{n},t_{n+1}],{\tilde{\ell}}_{ni}(t)={\tilde{\ell}}_{ni}(t_{n}+k_{n}\tau)={\tilde{\ell}}_{i}(\tau),i=0,\dotsc,q.$ Obviously, ~${\tilde{\ell}}_{ni}(\tilde{t}_{nj})=\delta_{ij},i,j=0,1,\dotsc,q,$ with $\tilde{t}_{nj}:=t_{n}+k_{n}\tilde{c}_{j},j=0,\dotsc,q.$ With this notation, let the interpolant $\widehat{v}$ of a time dependent function $v$ be the piecewise polynomial function approximating $v$ defined by interpolating the nodal values $v(\tilde{t}_{nj})$ of $v$ as follows
@@
$\begin{split}\mathcal{G}_{RK}(v)&=\|\widehat{v}_{t}(t)+I_{q-1}L\widehat{v}(t)-I_{q-1}f(t)\|^{2}_{L^{2}((0,T);L^{2}(\varOmega))}+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}\\ &=\int_{0}^{T}\|\widehat{v}_{t}(t)+I_{q-1}L\widehat{v}(t)-I_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}.\end{split}$ (26)
@@ 2.4.3 A general discrete Loss
of a time dependent function $v$ to be the piecewise polynomial function of local degree $q$ approximating $v.$ We assume that if $v$ is continuous in time, then $\widehat{v}$ is a piecewise polynomial globally continuous in time function. The standard interpolation operators as defined in Section 2.4.2 are typical cases of these operators. We are ready to define the time-discrete loss by
$\mathcal{G}_{k}(v)=\int_{0}^{T}\|\widehat{v}_{t}(t)+\varPi_{q-1}L\widehat{v}(t)-{\widetilde{\varPi}}_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}.\$ (28)
In Section 4 we show systematically that all Runge–Kutta methods considered and both continuous and discontinuous Galerkin time discrete schemes are associated to the general formulation involving the discrete operators $\varPi_{q-1},{\widetilde{\varPi}}_{q-1}$ and ${\widehat{\varPi}}_{q}.$
@@ 3 Stability and Convergence for Parabolic equations
{$&u_{t}+Lu=f\qquad&&\text{in }&&\varOmega_{T},\\ &u=0\qquad&&\text{on }&&\partial\varOmega\times(0,T],\\ &u=u_{0}\qquad&&\textrm{in }&&\varOmega\times\{t=0\}.\end{aligned}\right.$ (30)
@@
$\mathcal{G}(v)=\int_{0}^{T}\|v_{t}(t)+Lv(t)-f(t)\|^{2}_{L^{2}(\varOmega)}{\mathrm{d}}t+|v(0)-u_{0}|^{2}_{H^{1}(\varOmega)}\,.$ (31)
@@ 3.1 Time discrete training through Runge–Kutta and time Galerkin
$\mathcal{G}_{k}(v)=\int_{0}^{T}\|\widehat{v}_{t}(t)+\varPi_{q-1}L\widehat{v}(t)-{\widetilde{\varPi}}_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t+|\widehat{v}(0)-u_{0}|_{H^{1}(\varOmega)}^{2}.\$
@@
The neural network spaces are selected to meet specific approximability criteria aligned with established results in approximation theory; see, e.g., [Grohs_space_time_approx:2022, Mishra:appr:rough:2022, Xu] and their references. However, subtle challenges arise, particularly with boundary conditions and the fact that existing approximation results do not yet offer concrete guidance on selecting specific architectures. Nevertheless, to investigate the potential convergence of minimizers, we assume that the following approximability requirements are satisfied. The required smoothness of the spaces is guaranteed by selecting smooth enough activation functions. The neural network spaces are selected such that for each $\ell\in\mathbb{N}$ we associate a space $V_{\mathcal{N},0},$ which is denoted by $V_{\ell}$ with the approximation property: For each $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega))$ there exists a $w_{\ell}\in V_{\ell}$ such that
@@ 3.1.1 Stability-Maximal Regularity
Following [GGM_pinn_2023], we call our methods stable if two key properties, roughly stated as follows, hold
@@
Proof  Since
$\int_{0}^{T}\|\widehat{U}_{t}+L\overline{U}-{\widetilde{\varPi}}_{q-1}f(t)\|^{2}_{L^{2}(\varOmega)}\,{\mathrm{d}}t\,\leqslant C,$ (39)
@@ 3.2 Convergence of the minimizers
$\begin{split}\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell},\varphi^{\prime}{\rangle}{\mathrm{d}}t-&\int_{0}^{T}{\langle}\overline{U}_{\ell},\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}\leqslant\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},\varphi^{\prime}-I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}+\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}\\ &=\Big{|}\int_{0}^{T}{\langle}\widehat{U}_{\ell}-\overline{U}_{\ell},\varphi^{\prime}-I_{0}\varphi^{\prime}{\rangle}{\mathrm{d}}t\Big{|}.\end{split}$ (44)
@@
Proof  For $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega)\,)$, we choose a smooth approximation $(w_{\delta})\subset C^{\infty}([0,T];H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega)\,)$ such that
@@
Let $\mathcal{G},\;\mathcal{G}_{\ell}$ be the energy functionals defined in (31) and (14), respectively. Let $u$ be the exact solution of (30) and let $(u_{\ell}),$ $u_{\ell}\in V_{\ell},$ be a sequence of minimizers of $\mathcal{G}_{\ell}$, i.e.,
@@
Applying the Aubin–Lions Lemma, we conclude the existence of $\tilde{u}\in L^{2}((0,T);H^{1}(\varOmega))$ such that $\widehat{u}_{\ell}\rightarrow\tilde{u}$ in $L^{2}((0,T);H^{1}(\varOmega))$ up to a subsequence not re-labeled. Furthermore, the arguments in Lemma 3 show that $L\tilde{u}\in L^{2}((0,T);L^{2}(\varOmega))\,.$ Next we show that $\tilde{u}$ is the minimizer of $\mathcal{G},$ and hence $\tilde{u}=u.$ Indeed, let $w\in H^{1}((0,T);L^{2}(\varOmega))\cap L^{2}((0,T);H^{2}(\varOmega)\cap H^{1}_{0}(\varOmega))$, and let $(w_{\ell})$ be such that $w_{\ell}\to w$ and